OpenStudy (anonymous):

The figures below show two different ways of arranging four identical triangles of grey poster board on top of a white square. The square has sides equal to x + y, while the hypotenuse of each triangle is represented by the variable c. Hazel wrote the following statements to prove that c2 = x2 + y2. Which is the first incorrect statement in Hazel’s proof? Answer Statement 4 Statement 5 Statement 6 Statement 7

OpenStudy (anonymous):

Hazel wrote the following statements to prove that c2 = x2 + y2. 1. Area of the four grey triangles inside figure A = A equals four times the quantity one-half times x times y, which equals 2 times x times y. 2. Area of the white square inside figure A = c2 3. Area of figure A = c2 + 2xy 4. Area of the four grey triangles inside figure B = A equals four times the quantity one-half times x times y, which equals 2 times x times y. 5. Area of the two white squares inside figure B = x2+ y2 6. Area of figure B = 2x2+ 2y2 + 2xy 7. Area of figure A = area of Figure B, hence c2 + 2xy = 2x2+ 2y2 + 2xy 8. Therefore, c2 = x2+ y2

OpenStudy (anonymous):

i think it is statement 5 but im not positive of my answer... help please?

OpenStudy (anonymous):

statement 2 , how it will be a sqr? it is a rhombus. angle can not be 90 deg.

OpenStudy (anonymous):

statement 2 isnt one of my options though only statements 4 through 7 are.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yeahh..

OpenStudy (anonymous):

6, bcz, Total area of B=x^2+y^2+2xy

OpenStudy (anonymous):